OpenStudy (anonymous):

what is the simplified form of the quantity x-squared minus 36 over the quantity x minus 6? Select one: A. x - 6, with the restriction x ≠ -6 B. x - 6, with the restriction x ≠ 6 C. x + 6, with the restriction x ≠ 6 D. x + 6, with the restriction x ≠ -6

OpenStudy (anonymous):

\[\frac{ x^{2}-36 }{ x-6 }\]

OpenStudy (anonymous):

factorize the top x^2 - 36 = (x+6)*(x-6)

OpenStudy (anonymous):

\[\frac{ (x+6)*(x-6) }{ x-6 }\]

OpenStudy (anonymous):

the (x-6) cancel out, leaving you with just x+6

OpenStudy (anonymous):

but x cannot equal 6, because if it did, the denominator of the original equation would equal 0, and we cannot divide by 0

OpenStudy (anonymous):

so your answer is C

OpenStudy (anonymous):

Thank you! I now understand what the question means by "restriction." since x cannot equal 6, 6 is the restriction. Am I right? :P

OpenStudy (anonymous):

yes!